OpenStudy (anonymous):

Fill in the blank that completes the square: x^2+4/3x+ _____

OpenStudy (anonymous):

\[x ^{2}+\frac{ 4 }{ 3 }x+ blank\]

OpenStudy (anonymous):

\[(a+b)^2=a^2+2ab+b^2\] a=x b= ___